Solve the following. Show your solution and box your final answer.

Calculate the pH of a solution with an OH- ion concentration of 5.1x10-3 M.
1
Expert's answer
2020-11-26T05:10:16-0500

pOH=-log[OH-]

      =-log (5.1×10^-3)

      =2.292

pH=14-pOH

    =14-2.292

    pH=11.708